a dog is walking to the right at 1.5 (m/s) spies a cat ahead, and begins chasing the cat with a constant acceleration of 12 (m/s2). what is the velocity of the dog after runniing for 3.0 m?
1
Expert's answer
2019-10-02T09:51:34-0400

from the formula

"v^2-v_0^2=2\cdot a \cdot s"

Where

"v_0=1.5 m\/s" -starting speed

"a=15m\/s^2" -acceleration

"s=3 m" -the distance the dog ran

define "v"

"v=\sqrt{2\cdot a\cdot s\cdot + v_0^2}=\sqrt{2\cdot 12\cdot 3\cdot + 1.5^2}=
8.617 m/s"